# Blockwise SFT for Diffusion Language Models: Reconciling Bidirectional Attention and Autoregressive Decoding

### v1

@@ Abstract
Discrete diffusion language models have shown strong potential for text generation, yet standard supervised fine-tuning (SFT) misaligns with their semi-autoregressive inference: training randomly masks tokens across the entire response, while inference generates fixed-size blocks sequentially. This mismatch introduces noisy prefixes and leaky suffixes, biasing gradients away from the desired blockwise likelihood. We propose Blockwise SFT, which partitions responses into fixed-size blocks, selects one active block per step for stochastic masking, freezes all preceding tokens, and fully hides future ones. Loss is computed only over the active block, directly mirroring the blockwise decoding process. Experiments on GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), and MetaMathQA (Yu et al., 2024) show consistent gains over classical SFT under equal compute or token budgets. Block size consistency studies and ablations confirm that improvements stem from faithful training–inference alignment rather than incidental masking effects. Our results highlight the importance of matching supervision granularity to the decoding procedure in diffusion-based language models.
@@ 1 Introduction
Large language models based on discrete diffusion have moved from proofs of concept to competitive conditional generators (Gong et al., 2023). Intuitively, a diffusion LM learns to undo noise: during training it corrupts response tokens with a relaxed categorical process and predicts the clean tokens; during inference it starts from noise and iteratively denoises over multiple steps to produce text. In parallel, semi-autoregressive decoding has emerged as a practical middle ground between fully parallel refinement and strict left-to-right generation: systems such as SSD-LM (Han et al., 2023) generate fixed-size blocks sequentially, exploiting within-block parallelism while preserving cross-block causality. Modern dLLMs, including the LLaDA family (Nie et al., 2025), expose blockwise inference as a first-class interface. This setting raises a simple question that motivates our work: can a training objective based on bidirectional attention and full-sequence random masking serve a model that will be decoded causally, block by block?
@@
The prevailing supervised fine-tuning (SFT) recipe for diffusion LMs is misaligned with that use case. Classical SFT randomly masks tokens across the entire response and reconstructs them in a single pass. At inference, the model instead receives a clean, deterministic prefix and produces exactly one block while the future is fully hidden. As illustrated in Figure 1, this mismatch produces three concrete issues: (i) noisy prefixes—training conditions on corrupted contexts that never appear at test time; (ii) dependency leakage—training may reveal tokens inside or beyond the target block, violating the causal structure required at inference; and (iii) a granularity mismatch—loss is computed token by token while decisions at inference are made block by block.
Figure 1: Classical SFT, Blockwise SFT, and autoregressive inference.
We introduce Blockwise SFT to close this gap. The idea is to make training look like deployment. At each step, we focus learning on a single active block. The prefix before it is kept clean and frozen, the suffix after it is fully hidden, and the loss is computed only on the active block. This alignment preserves cross-block causality, concentrates supervision where the next decision is made, and is architecture-agnostic: any diffusion LM that decodes in blocks can adopt it without changing its network or sampler.
@@
The contributions of this work encompass three key aspects. We first diagnose why full-response SFT is ill-suited for semi-autoregressive, blockwise decoding in discrete diffusion LMs, showing that random masking creates noisy prefixes and dependency leakage (Figure 1). We then introduce Blockwise SFT, a drop-in objective that supervises one active block given a clean prefix and a hidden future, requiring no architectural changes or exotic schedules. Finally, we provide theory explaining when and why this alignment works: the objective forms a variational upper bound on blockwise likelihoods, simple block/time sampling yields unbiased gradient estimates, and the gradient bias of classical SFT under mismatch can be characterized precisely (see §3.3).
In short, aligning the training signal with the decoding procedure matters. Blockwise SFT turns a widely adopted inference pattern into an equally principled training objective, yielding models that are both more accurate and more efficient to fine-tune. We validate this claim by fine-tuning on MetaMathQA and evaluating on GSM8K and MATH, observing consistent gains under matched compute and matched supervision budgets, improved early-training efficiency, and state-of-the-art results in a head-to-head comparison (Table 1). Analyses further show that accuracy peaks when training and inference use the same block size (§4.5) and degrades with injected prefix noise or suffix visibility (§4.6).
@@ 2 Related Work
Discrete Diffusion Models for Text Generation
Diffusion has been extended to discrete text via Argmax Flows and Multinomial Diffusion (Hoogeboom et al., 2021), which map tokens to continuous space and add categorical noise during training. D3PMs (Austin et al., 2023) generalize this with structured corruption processes to better capture token relationships. DiffuSeq (Gong et al., 2023) applies diffusion to conditional sequence-to-sequence tasks, matching or surpassing autoregressive baselines. Recent evaluations (Weligalle, 2025) highlight trade-offs between modeling quality and generation speed.
Supervised Fine-Tuning Methods
For autoregressive LLMs, SFT trains on cross-entropy loss over response tokens (Ouyang et al., 2022), with variants such as instruction tuning (Wei et al., 2023; Chung et al., 2022) improving generalization. Diffusion-based LLMs use iterative refinement, with SFT approaches (Austin et al., 2023; Li et al., 2022) reconstructing randomly masked tokens in one step. This departs from the blockwise, semi-autoregressive decoding common in diffusion models, creating a training–inference mismatch.
Autoregressive Decoding in Diffusion-Based Language Models
Early discrete diffusion generators were slow and sometimes incoherent. SSD-LM (Han et al., 2023) introduced semi-autoregressive decoding: generating fixed-size blocks with diffusion denoising while freezing prior blocks. Block Diffusion (Arriola et al., 2025) interpolates between pure diffusion and fully autoregressive sampling via block decomposition and KV-caching, improving flexibility and throughput. Adaptive Parallel Decoding (APD) (Israel et al., 2025) adjusts block size by uncertainty, combining diffusion marginals with an autoregressive verifier to approach AR speeds with minimal quality loss.
Contemporary diffusion–SFT variants.
Beyond classical SFT, several recent variants target efficiency or stability rather than decoding alignment. MDLM introduces a simplified, substitution-based objective that reduces to a mixture of masked-LM losses and, with a strong training recipe, narrows the gap to autoregressive models (Sahoo et al., 2024). Soft-Masked Diffusion LM replaces hard corruption with linguistically informed soft-masking and applies per-step cross-entropy to better handle discrete tokens and reduce cost (Chen et al., 2023). RDM reparameterizes discrete diffusion, producing alternative training and sampling procedures that improve text generation quality (Zheng et al., 2024). Two-Step Loss & Scheduling explicitly mitigates train–inference discrepancy by training with a two-step diffusion and gradually increasing self-conditioning probability (Asada & Miwa, 2025). These approaches primarily refine the diffusion objective, noise schedule, or parameterization while continuing to supervise randomly across the whole response. In contrast, our Blockwise SFT aligns the granularity of supervision with semi-autoregressive, blockwise decoding (clean prefix, hidden future, loss on the active block). Under equal compute budget, this alignment yields consistent gains over all above variants on GSM8K and MATH (Table 1).
@@ 3 Methodology
3.1 Overview and intuition.
Classical SFT teaches a diffusion LM to fill in randomly masked tokens anywhere in the response, using bidirectional attention over whatever remains visible. Deployment looks different: semi-autoregressive decoding commits one block at a time, conditioning on a clean prefix and never peeking at future tokens. Imagine a short math line “x=53=x=5-3=\,?”. Training might corrupt the “5” or even show the answer token; inference never sees either event—it always receives an untouched prefix and must decide the next block causally.
Our remedy is to make training look like deployment. At each step we choose the active block—the next contiguous chunk the model would emit—keep the prefix exactly as it will appear at inference, fully hide the suffix, and place the loss only on the active block. The rest of this section turns this story into equations, states the three guarantees we rely on (one variational bound and two properties of the stochastic estimator), and defers proofs and estimator details to the appendix.
3.2 Classical Supervised Fine-Tuning for Discrete Diffusion and Its Limitations
Method recap.
Given an instruction–response sequence $\mathbf{x}_{1:L}=[\mathbf{c};\mathbf{r}]$, Classical SFT draws a binary mask $\mathbf{m}_{1:L}$ over response tokens, independently per token with probability $\pi$:
$m_{i}\sim\mathrm{Bernoulli}(\pi),\quad i\in\{L_{c}{+}1,\dots,L\}.$
Let $\tilde{\mathbf{x}}=\mathbf{x}\odot(1-\mathbf{m})$ be the masked input. With $T$ diffusion steps, forward kernel $q_{t}(\mathbf{z}_{t}\mid\mathbf{x})$, reverse model $\mathbf{p}_{\theta}(\mathbf{x}\mid\mathbf{z}_{t},t)$, and weights $\{\omega_{t}\}_{t=1}^{T}$, the loss is
@@ 3.2 Classical Supervised Fine-Tuning for Discrete Diffusion and Its Limitations
$\mathcal{L}_{\mathrm{SFT}}(\theta)$ $=\sum_{t=1}^{T}\omega_{t}\,\mathbb{E}_{\mathbf{x},\mathbf{m},\mathbf{z}_{t}}\Bigl{[}-\!\sum_{i=1}^{L}\!\mathbf{1}[m_{i}=1]\,\log p_{\theta}\bigl{(}x_{i}\mid\mathbf{z}_{t},t\bigr{)}\Bigr{]},$ (1)
i.e., an MLM-style cross-entropy at each diffusion step. Intuitively, the model learns to undo noise: training recovers clean tokens from a noised input; inference starts from noise and denoises iteratively to produce text.
Limitations under blockwise decoding.
Blockwise decoding commits decisions one block at a time. Let $B$ be the block size; index the next-to-generate block by $a$; write its token indices as $\mathcal{I}_{a}$; and let $\mathcal{I}_{\mathrm{prefix}}^{(a)}$ / $\mathcal{I}_{\mathrm{suffix}}^{(a)}$ be indices before / after block $a$. Relative to the blockwise risk in equation 3 (defined below), Classical SFT creates three specific mismatches:
(i) Prefix corruption. Training corrupts the conditioning context with probability
$\Pr\!\big{[}\exists\,i\in\mathcal{I}_{\mathrm{prefix}}^{(a)}:\,m_{i}=1\big{]}=1-(1-\pi)^{|\mathcal{I}_{\mathrm{prefix}}^{(a)}|},$
whereas inference always uses a clean, deterministic prefix.
(ii) Suffix leakage. Training can leave future tokens visible with probability
$\Pr\!\big{[}\exists\,j\in\mathcal{I}_{\mathrm{suffix}}^{(a)}:\,m_{j}=0\big{]}=1-\pi^{|\mathcal{I}_{\mathrm{suffix}}^{(a)}|},$
allowing bidirectional attention to exploit information that will be hidden at deployment.
(iii) Granularity mismatch. The loss in equation 1 spreads supervision across all response positions. In expectation, the supervised mass on the active block is $\pi\,|\mathcal{I}_{a}|$, identical (up to size) to that allocated to non-active blocks. In contrast, decoding cares about the next block, which we formalize next; the dispersion of learning signal leads to the bias quantified by Theorem 3.3.
3.3 Deriving the Blockwise SFT Objective
From decoding to a training risk.
@@ 3.3 Deriving the Blockwise SFT Objective
Write the response $\mathbf{r}$ as $M$ consecutive blocks $\mathbf{b}^{(1)},\dots,\mathbf{b}^{(M)}$ after the instruction $\mathbf{c}$. Under semi-autoregressive decoding, the sequence likelihood factorizes as
@@
where $\mathbf{c\!ontext}^{(a)}=\mathbf{x}_{1:L_{c}+B(a-1)}$ is the clean prefix up to block a1a{-}1, and $\mathbf{b}^{(a)}=\mathbf{x}_{L_{c}+B(a-1)+1:L_{c}+Ba}$ are tokens of block $a$. The risk we care about is the sum of blockwise NLLs,
$\mathcal{R}_{\text{block}}(\theta)=\mathbb{E}_{\mathbf{x}}\!\left[-\sum_{a=1}^{M}\log p_{\theta}\!\bigl{(}\mathbf{b}^{(a)}\mid\mathbf{c\!ontext}^{(a)},t{=}0\bigr{)}\right],$ (3)
so training should align its gradients with equation 3, not with a full-response denoising objective that mixes prefixes and leaks suffix information.
A block-local diffusion surrogate.
To match decoding, we select an active block $a$, keep the prefix clean (no gradient), fully hide the suffix, and apply diffusion cross-entropy only within $a$. Let $\mathcal{I}_{a}$ be indices of block $a$, and let $\mathbf{z}_{t}\!\sim\!q_{t}(\cdot\mid\mathbf{x})$ be the noised state at step $t$. Define
$\tilde{\mathcal{L}}_{t}(\theta;\mathbf{x},a)=-\!\!\sum_{i\in\mathcal{I}_{a}}\log p_{\theta}\!\bigl{(}x_{i}\mid\mathbf{z}_{t},t\bigr{)},\qquad\mathcal{L}_{\text{BW--SFT}}(\theta)=\mathbb{E}_{\mathbf{x},\,a}\!\Bigl{[}\sum_{t=1}^{T}\omega_{t}\,\mathbb{E}_{\mathbf{z}_{t}}\bigl{[}\tilde{\mathcal{L}}_{t}(\theta;\mathbf{x},a)\bigr{]}\Bigr{]}.$ (4)
Single-$t$ sampling and importance weighting that recover the $t$-sum in equation 4 are given in Appendix A.5.
Why this surrogate is faithful.
Applying the discrete-diffusion ELBO only to the masked active block—conditioning on the clean prefix and hiding the suffix—yields a variational upper bound on the blockwise NLL.
Theorem 3.1 (Variational upper bound for blockwise likelihood).
For suitable nonnegative weights $\{\omega_{t}\}_{t=1}^{T}$ and a constant $C$ independent of $\theta$,
@@
and hence $\mathcal{R}_{\text{block}}(\theta)\leq\mathcal{L}_{\text{BW--SFT}}(\theta)+C^{\prime}$ with $C^{\prime}$ independent of $\theta$. Proof in Appendix A.6.
Unbiased gradients under block/time sampling.
Sampling the active block and the diffusion step yields an unbiased estimator of the surrogate gradient; appropriate reweighting recovers the full objective.
Theorem 3.2 (Unbiasedness of blockwise gradients; importance weighting).
Let $a\sim\rho$ over $\{1,\dots,M\}$ and $t\sim\tilde{\omega}$ with $\tilde{\omega}_{t}=\omega_{t}/\sum_{s=1}^{T}\omega_{s}$. Define
$\widehat{g}(\theta)=\frac{1}{\rho(a)}\!\left(\sum_{s=1}^{T}\omega_{s}\right)\nabla_{\theta}\,\mathbb{E}_{\mathbf{z}_{t}\sim q_{t}(\cdot\mid\mathbf{x})}\bigl{[}\tilde{\mathcal{L}}_{t}(\theta;\mathbf{x},a)\bigr{]}.$
Then $\mathbb{E}[\widehat{g}(\theta)]=\nabla_{\theta}\,\mathcal{L}_{\text{BW--SFT}}(\theta)$ for any $\rho$. If a local fidelity condition holds, $\mathbb{E}[\widehat{g}(\theta)]=\nabla_{\theta}\,\mathcal{R}_{\text{block}}(\theta)$ (up to a global scalar). Proof in Appendix A.6.
Why classical SFT is biased under mismatch.
Noisy prefixes and leaky suffixes induce a systematic gradient bias when training with Classical SFT but decoding blockwise.
Theorem 3.3 (Classical SFT gradient bias under mismatch).
Let $\nabla_{\theta}\ell^{\star}(\theta;\mathbf{x},a)$ be the clean active-block gradient $\nabla_{\theta}[-\log p_{\theta}(\mathbf{b}^{(a)}\mid\mathbf{c\!ontext}^{(a)},t{=}0)]$, and $\nabla_{\theta}\ell^{\text{cls}}(\theta;\mathbf{x},a)$ be the Classical-SFT gradient with mask rate $\pi$. If the gradient map is $L_{\text{pre}}$-Lipschitz w.r.t. prefix perturbations and $L_{\text{suf}}$-Lipschitz w.r.t. suffix perturbations (holding the active block fixed), then
Proof in Appendix A.6.
Takeaways.
The surrogate in equation 4 (i) upper-bounds the blockwise NLL (Theorem 3.1), (ii) admits unbiased stochastic gradients via block/time sampling (Theorem 3.2), and (iii) corrects the gradient bias inherent to Classical SFT under blockwise decoding (Theorem 3.3). This explains why aligning the objective with the decoding procedure is both principled and effective.
3.4 Practical Implementation of Blockwise SFT
Objective.
Let $\mathcal{I}_{a}$ be the indices of the active block $a$. The Blockwise SFT loss is
$\mathcal{L}_{\mathrm{BW\!-\!SFT}}(\theta)=\sum_{t=1}^{T}\omega_{t}\;\mathbb{E}_{\mathbf{x},\,a,\,\mathbf{z}_{t}}\Bigl{[}-\sum_{i\in\mathcal{I}_{a}}\log p_{\theta}(x_{i}\mid\mathbf{z}_{t},t)\Bigr{]},$ (7)
with gradients back-propagated only through $\mathcal{I}_{a}$, matching the semi-autoregressive factorization in §3.3.
Algorithm.
Algorithm 1 gives a single training step for Blockwise SFT. In brief: partition the response into fixed-size blocks, sample one active block, keep the prefix clean and the suffix hidden, sample a mask rate and a diffusion step, compute the block-local diffusion loss, and update $\theta$ using gradients from that block only (notation matches §3.3).

### v2

@@ Abstract
Discrete diffusion language models have shown strong potential for text generation, yet standard supervised fine-tuning (SFT) misaligns with their semi-autoregressive inference: training randomly masks tokens across the entire response, while inference generates fixed-size blocks sequentially. This mismatch introduces noisy prefixes and leaky suffixes, biasing gradients away from the desired blockwise likelihood. We propose Blockwise SFT, which partitions responses into fixed-size blocks, selects one active block per step for stochastic masking, freezes all preceding tokens, and fully hides future ones. Loss is computed only over the active block, directly mirroring the blockwise decoding process. Experiments on GSM8K, MATH, and MetaMathQA show consistent gains over classical SFT under equal compute or token budgets. Block size consistency studies and ablations confirm that improvements stem from faithful training–inference alignment rather than incidental masking effects. Our results highlight the importance of matching supervision granularity to the decoding procedure in diffusion-based language models.
@@ 1 Introduction
Discrete diffusion models have emerged as a promising alternative to autoregressive language models for text generation. Unlike traditional left-to-right decoding, diffusion language models (LMs) learn to iteratively denoise corrupted text by predicting clean tokens from noisy versions (Gong et al., 2023). During training, these models corrupt response tokens with a relaxed categorical process and learn to recover the original text; during inference, they start from pure noise and progressively refine it into coherent text through multiple denoising steps.
Recent advances have explored semi-autoregressive decoding as a practical compromise between fully parallel generation and strict sequential processing. Systems like SSD-LM generate text in fixed-size blocks, where tokens within each block are produced in parallel while maintaining causal dependencies across blocks (Han et al., 2023). Modern discrete diffusion LMs, including the LLaDA family, have adopted blockwise inference as their primary generation strategy (Nie et al., 2025). This widespread adoption of blockwise decoding raises a fundamental question: can a training objective based on bidirectional attention and full-sequence random masking effectively serve a model that will be decoded causally, block by block?
@@
The prevailing supervised fine-tuning (SFT) approach for diffusion LMs is fundamentally misaligned with blockwise inference. Classical SFT randomly masks tokens across the entire response and reconstructs them using bidirectional context in a single pass. However, during inference, the model operates differently: it receives a clean, deterministic prefix and must produce exactly one block while all future tokens remain completely hidden. As illustrated in figure 1, this training-inference mismatch creates three critical problems: (i) noisy prefixes: the model trains on corrupted contexts that never appear during generation; (ii) dependency leakage: training may reveal tokens within or beyond the target block, violating the causal constraints enforced at inference; and (iii) granularity mismatch: training optimizes token-level decisions while inference requires block-level coordination.
Figure 1: Mismatch between training and inference: Classical SFT uses bidirectional attention and full-sequence masking, exposing future tokens and corrupting parts of the prefix; semi-autoregressive decoding is one-directional, conditioning only on a clean prefix while future blocks are strictly hidden.
To address this misalignment, we introduce Blockwise SFT (Figure 2), a training objective that mirrors deployment: at each step we supervise a single active block, keep the prefix clean and frozen, hide all future tokens, and compute loss only within the active block. This aligns gradients with blockwise decoding, preserves cross-block causality, and concentrates supervision where decisions are actually made. The recipe is architecture-agnostic and applies to any diffusion LM that supports blockwise inference.
Figure 2: Blockwise SFT step: split the completion, sample an active block, keep a clean prefix and a hidden future, and train only on masked tokens inside that block.
@@
Contributions. (1) We diagnose the core mismatch between classical full-response SFT and blockwise semi-autoregressive decoding, showing how global random masking induces noisy prefixes and suffix leakage. (2) We propose Blockwise SFT, a drop-in objective that supervises one active block at a time under a clean prefix and hidden future, requiring no architectural or sampling changes. (3) We provide theory: a variational upper bound on blockwise likelihoods, unbiased timestep-sampled gradients, and a characterization of the gradient bias in classical SFT under mismatch.
Aligning training with inference yields concrete gains. Fine-tuning on MetaMathQA and evaluating on GSM8K and MATH, Blockwise SFT consistently improves Pass@1 under matched FLOPs and matched supervised tokens, with faster early-training progress. Performance peaks when the training block size matches the inference block size and drops with injected prefix noise or suffix visibility, corroborating our alignment analysis.
@@ 2 Related Work
Discrete Diffusion Models for Text Generation Diffusion has been extended to discrete text via Argmax Flows and Multinomial Diffusion, which map tokens to continuous space and add categorical noise during training (Hoogeboom et al., 2021). D3PMs generalize this with structured corruption processes to better capture token relationships (Austin et al., 2023). DiffuSeq applies diffusion to conditional sequence-to-sequence tasks, matching or surpassing autoregressive baselines (Gong et al., 2023). Recent evaluations highlight trade-offs between modeling quality and generation speed (Weligalle, 2025).
Supervised Fine-Tuning Methods For autoregressive models, SFT trains on cross-entropy loss over response tokens (Ouyang et al., 2022), with variants such as instruction tuning improving generalization (Wei et al., 2023; Chung et al., 2022). Diffusion-based models use iterative refinement, with SFT approaches reconstructing randomly masked tokens in one step (Austin et al., 2023; Li et al., 2022). This departs from the blockwise, semi-autoregressive decoding common in diffusion LLMs, creating a training–inference mismatch.
Autoregressive Decoding in Diffusion-Based Language Models Early discrete diffusion generators were slow and sometimes incoherent. SSD-LM introduced semi-autoregressive decoding: generating fixed-size blocks with diffusion denoising while freezing prior blocks (Han et al., 2023). Block Diffusion interpolates between pure diffusion and fully autoregressive sampling via block decomposition and KV-caching, improving flexibility and throughput (Arriola et al., 2025). Adaptive Parallel Decoding (APD) adjusts block size by uncertainty, combining diffusion marginals with an autoregressive verifier to approach AR speeds with minimal quality loss (Israel et al., 2025).
Contemporary diffusion–SFT variants. Recent variants emphasize efficiency or stability rather than decoding alignment: MDLM mixes masked-LM–style losses to narrow the gap to autoregressive models (Sahoo et al., 2024); Soft-Masked Diffusion LM uses linguistically informed soft masking with per-step cross-entropy to reduce cost (Chen et al., 2023); RDM reparameterizes the discrete process to improve training and sampling (Zheng et al., 2024); and Two-Step Loss & Scheduling adopts a two-step diffusion with gradually increased self-conditioning to mitigate mismatch (Asada & Miwa, 2025). These methods chiefly modify the objective, schedule, or parameterization while supervising across the full response. In contrast, Blockwise SFT aligns supervision with semi-autoregressive blockwise decoding (clean prefix, hidden future, loss on the active block), yielding consistent gains on GSM8K and MATH under equal compute (Table 1).
@@ 3 Methodology
3.1 Overview: From Training-Inference Mismatch to Blockwise Alignment
Semi-autoregressive diffusion models face a fundamental challenge: they are trained to denoise tokens anywhere in a sequence, but deployed to generate specific blocks conditioned on clean prefixes. This section develops Blockwise SFT to bridge this gap.
Consider generating a simple expression like “53=5-3=\,?”. During training, classical methods might mask the “5” or reveal the answer “2” in the suffix, but scenarios that never occur during inference, which always starts from a pristine prefix and generates the next block without future information. This mismatch creates three specific issues we quantify in Section 3.2: (i) training corrupts prefixes that inference keeps clean, (ii) training leaks future tokens that inference never sees, and (iii) training spreads supervision across all positions while inference focuses on the next block.
Our solution: blockwise alignment. We make training mirror deployment. At each step, we identify the active block, which is the next chunk the model would generate during inference. We keep its prefix exactly as it appears at deployment (clean and fixed), completely hide the suffix (no future leakage), and concentrate the training signal solely on this active block. This alignment ensures that gradient updates directly optimize the objective used during generation.
Roadmap. Section 3.2 analyzes classical SFT’s limitations under blockwise decoding and quantifies the resulting gradient bias. Section 3.3 derives our blockwise objective from first principles, proving it provides an unbiased variational bound on the true decoding risk. Section 3.4 translates the theory into a practical algorithm. Throughout, we interleave formal results with intuitive explanations to maintain clarity.
3.2 Classical SFT: Analysis of Training-Inference Mismatch
We first recall the classical approach, then characterize how it diverges from blockwise decoding.
Classical SFT recap. Given an instruction–response sequence $\mathbf{x}_{1:L}=[\mathbf{c};\mathbf{r}]$, classical SFT samples independent masks on response tokens with rate $\pi$ (i.e., $m_{i}\!\sim\!\mathrm{Bernoulli}(\pi)$ for $i\!>\!L_{c}$) and trains the model to reconstruct only the masked positions at each diffusion step $t$:
@@ 3.2 Classical SFT: Analysis of Training-Inference Mismatch
$\mathcal{L}_{\mathrm{SFT}}(\theta)=\sum_{t=1}^{T}\omega_{t}\,\mathbb{E}_{\mathbf{x},\mathbf{m},\mathbf{z}_{t}}\Bigl[-\!\!\sum_{i>L_{c}}\!\mathbf{1}[m_{i}=1]\,\log p_{\theta}\bigl(x_{i}\mid\mathbf{z}_{t},t\bigr)\Bigr].$ (1)
This objective treats masked positions uniformly and leverages bidirectional context—a capability unavailable during blockwise generation.
Quantifying the mismatch. At inference, the model generates block $a$ (indices $\mathcal{I}_{a}$) conditioned on a clean prefix $\mathcal{I}_{\mathrm{prefix}}^{(a)}$ and without access to the future $\mathcal{I}_{\mathrm{suffix}}^{(a)}$. Classical training violates all three conditions:
  • Prefix corruption. Random masking almost surely corrupts long prefixes (e.g., with $\pi{=}0.3$ and 100 prefix tokens, corruption occurs with probability $1-3.3\times 10^{-16}$), whereas inference always conditions on a pristine prefix.
  • Suffix leakage. A nontrivial fraction of future tokens remains visible during training even for modest suffix lengths, enabling shortcuts that are disallowed at inference.
  • Diluted supervision. The loss in Eq. equation 1 spreads gradients over all response positions. The active block obtains only a $|\mathcal{I}_{a}|/L_{r}$ share of the total signal, which is identical to any other block despite being the sole target at inference.
These mismatches induce a systematic bias in the learning signal:
Theorem 3.1 (Gradient bias under training-inference mismatch).
Let $\nabla_{\theta}\ell^{\star}(\theta;\mathbf{x},a)$ denote the ideal gradient for generating block $a$ from a clean prefix (as in inference), and $\nabla_{\theta}\ell^{\text{cls}}(\theta;\mathbf{x},a)$ the gradient from classical SFT. If gradients are $L_{\text{pre}}$-Lipschitz with respect to prefix perturbations and $L_{\text{suf}}$-Lipschitz with respect to suffix visibility, then:
Intuition. The bound quantifies how prefix corruption and suffix visibility pull gradients away from the true objective; the bias grows with sequence length and is not removed by tuning $\pi$, motivating a structurally aligned training procedure. (Proof in Appendix A.6.)
3.3 Deriving the Blockwise Objective
We now develop an objective that aligns training with inference, ensuring unbiased gradient estimates for blockwise generation.
@@ 3.3 Deriving the Blockwise Objective
Target: blockwise generation risk. Semi-autoregressive decoding generates a response as $M$ consecutive blocks $\mathbf{b}^{(1)},\dots,\mathbf{b}^{(M)}$ following instruction $\mathbf{c}$. The likelihood factorizes as:
@@
where $\mathbf{context}^{(a)}=\mathbf{x}_{1:L_{c}+B(a-1)}$ represents the clean prefix through block a1a{-}1. The risk we minimize is:
$\mathcal{R}_{\text{block}}(\theta)=\mathbb{E}_{\mathbf{x}}\!\left[-\sum_{a=1}^{M}\log p_{\theta}\!\bigl(\mathbf{b}^{(a)}\mid\mathbf{context}^{(a)},t{=}0\bigr)\right].$ (4)
This is the exact objective optimized during inference.
Constructing a faithful training surrogate. Direct optimization of Eq. equation 4 is intractable due to the discrete diffusion process. We derive a surrogate that maintains three critical properties: it upper-bounds the true risk, admits efficient stochastic gradient estimation, and preserves the blockwise structure.
For active block $a$ with indices $\mathcal{I}_{a}$, we apply diffusion training only within this block while keeping the prefix clean and suffix hidden:
$\tilde{\mathcal{L}}_{t}(\theta;\mathbf{x},a)=-\!\!\sum_{i\in\mathcal{I}_{a}}\log p_{\theta}\!\bigl(x_{i}\mid\mathbf{z}_{t},t\bigr),$ (5)
where $\mathbf{z}_{t}\!\sim\!q_{t}(\cdot\mid\mathbf{x})$ incorporates noise at diffusion step $t$. The full objective averages over blocks and diffusion steps:
$\mathcal{L}_{\text{BW--SFT}}(\theta)=\mathbb{E}_{\mathbf{x},\,a}\!\Bigl[\sum_{t=1}^{T}\omega_{t}\,\mathbb{E}_{\mathbf{z}_{t}}\bigl[\tilde{\mathcal{L}}_{t}(\theta;\mathbf{x},a)\bigr]\Bigr].$ (6)
This construction provides rigorous guarantees connecting training to inference:
Theorem 3.2 (Variational bound on blockwise generation).
For appropriate weights $\{\omega_{t}\}_{t=1}^{T}$, the blockwise diffusion objective upper-bounds the generation risk:
@@
where $C$ is independent of $\theta$. Therefore, $\mathcal{R}_{\text{block}}(\theta)\leq\mathcal{L}_{\text{BW--SFT}}(\theta)+C^{\prime}$.
Intuition. By focusing diffusion training on individual blocks with proper conditioning, we obtain a tractable upper bound on the exact inference objective. Minimizing this bound improves generation quality. (Proof in Appendix A.6.)
Theorem 3.3 (Unbiased gradient estimation).
Sampling block $a\sim\rho$ and diffusion step $t\sim\tilde{\omega}$ (where $\tilde{\omega}_{t}\propto\omega_{t}$) yields an unbiased gradient estimator:
$\widehat{g}(\theta)=\frac{1}{\rho(a)}\!\left(\sum_{s=1}^{T}\omega_{s}\right)\nabla_{\theta}\,\mathbb{E}_{\mathbf{z}_{t}\sim q_{t}(\cdot\mid\mathbf{x})}\bigl[\tilde{\mathcal{L}}_{t}(\theta;\mathbf{x},a)\bigr].$ (8)
Then $\mathbb{E}[\widehat{g}(\theta)]=\nabla_{\theta}\,\mathcal{L}_{\text{BW--SFT}}(\theta)$ for any sampling distribution $\rho$.
Intuition. Despite sampling only one block and one diffusion step per update, our gradients remain unbiased estimates of the full objective. This enables efficient training while maintaining theoretical soundness. (Proof in Appendix A.6.)
Summary. The blockwise objective in Eq. equation 6 bridges the gap between tractable training and faithful inference by: (i) providing a variational bound on the true generation risk, (ii) enabling unbiased stochastic optimization, and (iii) eliminating the systematic bias inherent in classical approaches.
3.4 Practical Implementation
We now translate the theoretical framework into a practical training algorithm. Algorithm 1 implements one training step of Blockwise SFT. The key insight is simplicity: we modify only the masking strategy, leaving the model architecture and inference procedure unchanged.